Lewis and Clark Camping Supplies Inc. is borrowing $54,000 from Western State Bank. The total interest is $20,800. The loan will be paid by making equal monthly payments for the next three years. What is the effective rate of interest on this installment loan?
Answer #1

Loan amount(PV) = $ 54000

Interest paid = $ 20,800

Number of years(nper) = 3 = 36 months

Total amount repaid = 54,000 + 20,800 = $ 74,800

Monthly payments on the loan (PMT) = 74,800/36 = $ 2077.77

Rate of interest = = 1.88% per month =

22.55% per annum